Strontium 90 is a radioactive material that decays according to the function
A(t)=A0e−0.0244t, where A0 is the initial amount present and A is the amount present at time t​ (in years). Assume that a scientist has a sample of
400 grams of strontium 90.
a) what is the decay rate of strontium 90?
b) how much strontium 90 is left after 40 years
c) when will only 300 grams of strontium 90 be left?
d) what is the half life of strontium 90?

Answer by Boreal(15235)

decay rate is 2.44% per year
after 40 years, 400*(e^-0.976)=150.73 gm
300=400(e^-0.0244t)
0.75=e^-0.0244t
ln both sides
-0.2877=-0.0244t
t=11.79 years
t (1/2) occurs when ln(0.5)=-0.0244t
t=28.41 years